OpenStudy (anonymous):

7^(x + 1) = 2^3x. Solve for x.

mathslover (mathslover):

\[\huge{7^{x+1}=2^3 x}\] \[\huge{\frac{7^{x+1}}{2^3}=x}\] \[\huge{\frac{7^{x+1}}{8}=x}\]

OpenStudy (anonymous):

I'm sure the OP meant 2 to the 3X power

OpenStudy (anonymous):

yeah, sorry, 2^(3x)

OpenStudy (callisto):

\[7^{(x + 1)} = 2^{3x} \]\[(x+1)log7 = 3xlog2\]\[log7 = 3xlog2-xlog7\]\[log7 = x(3log2-log7)\]\[ x = \frac{log7}{(3log2-log7)}\]

OpenStudy (anonymous):

I'm trying to wrap my head around that.

OpenStudy (anonymous):

I don't get the third step, log7=3xlog2−xlog7.

OpenStudy (anonymous):

How did you go from the second step to the third?

OpenStudy (anonymous):

That was correct, thanks!

OpenStudy (callisto):

Sorry, I skipped a step there.. (x+1)log7=3xlog2 xlog7 + log7 = 3xlog2 log7=3xlog2−xlog7.

OpenStudy (zepp):

Callisto distributed log7 to stuffs in parentheses